5. A solution of 10g of sodium chloride (NaCl) is added to 100g water (H2O) in an attempt to elevate the boiling point. What is the boiling point of the solution? kb of water is 0.51°C/m.

1
Expert's answer
2021-11-08T08:14:16-0500

"\u0394\nT\nb\n=\ni\n\u22c5\nK\nb\n\u22c5\nb"

10/58.44= 0.17mol

0.17/0.1= 1.7molal

1.7×0.51= 0.867

= 100+0.867= 100.87°C